OpenStudy (anonymous):

If g(x) = 2(x − 4), find the value of x if g(x) = 20.

OpenStudy (anonymous):

32

OpenStudy (anonymous):

20 = 2(x-4) 10 = x - 4 14 = x

OpenStudy (anonymous):

2(x-4) = 20 2x - 8 = 20 2x = 20 + 8 2x = 28 x = 28/2 x = 14

OpenStudy (anonymous):

annas solved g(20).

OpenStudy (savvy):

32

OpenStudy (anonymous):

Not the problem being asked.

OpenStudy (anonymous):

Savvy made the same mistake.

OpenStudy (anonymous):

Thank you (: This one always gets me confused.

OpenStudy (savvy):

yeah..!!!! i made it...

OpenStudy (anonymous):

@Rainbow.ninja you got this question earlier !! don't post them twice for no reason ! http://openstudy.com/study#/updates/4f7c9f56e4b09f22231b9110 there was nothing that any random calculator could do !

OpenStudy (anonymous):

Sorry ..>.>

OpenStudy (anonymous):

Once you understand that a function has a value for different values of its input, you will understand these problems rather easily. In this case, you are given the value of the function and being asked what "input" produces that value. In the other case, you are given the input and asked what value does the function have.

OpenStudy (anonymous):

@kreshnik She posted again because she never got a confirmation whether the answer was 14. She did ask if that was right.

OpenStudy (anonymous):

Thank you @gt

OpenStudy (anonymous):

when she asked I wasn't there... she didn't mentioned me so I could confirm. but anyway... my mother is a teacher.. she teaches kids like 7-8 years old. I'm SURE they could have solved the rest of the task I left for @Rainbow.ninja I think we should HELP not DO their homework @GT

OpenStudy (anonymous):

He was HELPING me..